OpenStudy (anonymous):

will medal for straight answer. Which is the best deal for a package of soap? A. $5.75 for 14 oz B. $5.50 for 13 oz C. $3.45 for 3.5 oz D. $5.25 for 12 oz

OpenStudy (anonymous):

To find the answer you divide the dollar amount by the number of ounces to find out how much you're paying per ounce. 5.75/14 = .41 5.50/13 = .42 3.45/3.5 = .99 5.25/12 = .44 .41 is the lowest amount, so therefore it's the best buy :) Hope that helped

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

No problem :)